Assume that the world population is 10 billion people and each one is assigned a serial number between 0000000000 and 9999999999. If you pick a random person and look at the number, what is the probability that 5 digits are equal to "5" or more?
Answer #1

P(one digit from 0 to 9 is 5 or more) = Number of digits that are 5 or more/Total number of digits

= 5/10

= 0.5

Number of digits in a serial number = 10

Binomial distribution: P(X) = nCx px qn-x

n = 10

p = 0.5

q = 1 - p = 0.5

P(5 digits are equal to 5 or more) = P(X = 5)

= 10C5 x 0.55 x 0.55

= 0.2461